40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 31 21:12:05 2013, manasa, Update, General, Pump down called off 
    Reply  Thu Aug 1 16:03:52 2013, Steve, Update, General, reopen BS & IOO IOOnorth.jpg
    Reply  Thu Aug 1 22:49:31 2013, manasa, Update, General, Green status after PR3 flipping investigated 
       Reply  Fri Aug 2 17:50:10 2013, Jenne, Update, General, PR3 wedge angle adjusted 
          Reply  Fri Aug 2 22:51:10 2013, Jenne, Update, General, All vent tasks complete, just need oplev check 
             Reply  Mon Aug 5 18:02:34 2013, manasa, Update, General, Close up list checked 
             Reply  Tue Aug 6 07:52:58 2013, manasa, Update, General, Ready to pump down MCspot_0806.pngOSEM_0806.pngOL_0806.pngQPD_0806.png
Message ID: 8965     Entry time: Mon Aug 5 18:02:34 2013     In reply to: 8962
Author: manasa 
Type: Update 
Category: General 
Subject: Close up list checked 
  •  
  • Center beam on all AS optics
  • GET CAMERA IMAGES OF EVERYTHING

    • We must get images right before closing, right after closing, etc.
  • Make sure REFL is clear
    • dither PRM, see motion on AP tables
  • Make sure AS is clear
    • dither BS/ITM, see motion on AP tables
  • Using IPANG/POS pick-off mirrors, center beams on:
    • IPPOS
    • IPANG (aligned low in pitch)
  • Check green alignment in the arms and make sure the transmitted green reaches the PSL table.
  • Check all OpLevs centered, in and out of vacuum

    [Jenne, Manasa]

    IPANG needed to be re-aligned today. Heavy doors are in place and bolts tight (torque 25 & 45).

    Steve! We are ready for pump down!

    I will check the IFO alignment once again early tomorrow morning before Steve starts pumping down.

     

     

     

     

     

ELOG V3.1.3-